OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. B(6, 2), m =-1/2

OpenStudy (anonymous):

he general form of line i.e,slope point form is: y−y′=m(x−x′) where \[(x \prime,y \prime)=(6,2)\] & m=−12 just put it get the answer